OpenStudy (anonymous):

Find the local maximum and minimum values of f using both the First and Second Derivative Tests. f(x)=x/(x^2+16)

OpenStudy (anonymous):

take first derivative, set it equal to 0, solve for x. also find where the first derivative is udnefined..

OpenStudy (anonymous):

find all of those x values and then take 2nd derivatives..

OpenStudy (anonymous):

plug ur values into the 2nd derivative and whenever u get negative then u have max, and positive = min

OpenStudy (anonymous):

+4 and -4

OpenStudy (anonymous):

No the derivative, not the actual thing

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{x}{x^2+16}=\frac{16-x^2}{(x+16)^2}\] zeros are at 4 and-4

OpenStudy (anonymous):

oh okay so she was right.. and als x = -16 too

OpenStudy (anonymous):

??

OpenStudy (anonymous):

oh no that is not right

OpenStudy (anonymous):

-4 is a min, 4 is a max sorry

OpenStudy (anonymous):

nothing like revisionist history

OpenStudy (anonymous):

do you really need second derivative test here? it is clear as day that \[(x+16)^2>0\] for all x, so only need sign of \[16-x^2\] which is obviously negative on \[(-\infty,-4)\cup(4,\infty)\] and positive on \[(-4,4)\] making -4 a local min and 4 a local max

OpenStudy (anonymous):

Satellite you missed x = -16 as a local extremum!

OpenStudy (anonymous):

but you can use it if you want. the second derivative is \[\frac{2x(x^2-48)}{(x+16)^2}\]

OpenStudy (anonymous):

Satellite -16 is another extremum u should test lol sorry if im annoying u!

OpenStudy (anonymous):

hold the phone. why is -16 an extremum? i mean i believe you i just don't see it?

OpenStudy (anonymous):

In the denominator: (x+16)^2 there are extrema when 1st derivative is 0 or when it doesn't exist

OpenStudy (anonymous):

At x = -16 denominator is 0, there's a cusp

OpenStudy (anonymous):

derivative is \[\frac{16-x^2}{(x^2+16)^2}\] yes?

OpenStudy (anonymous):

the denominator is always positive

OpenStudy (anonymous):

in fact it is always greater than 16

OpenStudy (anonymous):

ohh wait is it? eariler u said it was just x not x^2

OpenStudy (anonymous):

let me look at the problem again

OpenStudy (anonymous):

I didn't take the derivative, sorry, no i meant according to ur 2nd or something post it would've been that..

OpenStudy (anonymous):

oh damn. i made a typo twice sorry. it is \[\frac{d}{dx}\frac{x}{x^2+16}=\frac{16-x^2}{(x^2+16)^2}\]

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

i forgot the square on the x

OpenStudy (anonymous):

Oh okay, then you are good..

OpenStudy (anonymous):

and second derivative is \[\frac{2x(x^2-48)}{(x^2+16)^2}\]

OpenStudy (anonymous):

sorry that was my fault

OpenStudy (anonymous):

extrema are correct though

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so is the answer -1/8 and 1/8

OpenStudy (anonymous):

are you there?